Give an upper bound and a lower bound for the expression 1/(a^(4)+3a^(2)+1) if a∈R


1
Expert's answer
2021-01-28T05:01:13-0500

The function "f(a)=\dfrac{1}{a^4+3a^2+1}" satisfies "0<f(a)\leq1" for all "a\in\mathbb{R}". Moreover, "f(0)=1" and "f(a)" tends to 0 as "a" tends to "\pm\infty". Then the upper bound of "\dfrac{1}{a^4+3a^2+1}" is 1 and its lower bound is 0.